OpenStudy (anonymous):

Math check. PART A: The slope of f(x) is 10 because. Answer: By picking two point on the table. ( 0, -1 ) and ( 1, 9 ) We can now find the slope. 9- (-1) M ------------------ = 10 1- 0 The slope of g(x) is 5 because. The equation form y=mx+b has the graph of the line with a slope of m and the y-intercept b. This is why the slope is 5. PART B: Which function has a greater y-intercept? Answer: g(x) has a greater y-intercept because it is positive 1 and the y-intercept from the table is -1.

OpenStudy (anonymous):

PART A: The slope of f(x) is 10 because. Answer: By picking two point on the table. ( 0, -1 ) and ( 1, 9 ) We can now find the slope. 9- (-1) M ------------------ = 10 1- 0 The slope of g(x) is 5 because. The equation form y=mx+b has the graph of the line with a slope of m and the y-intercept b. This is why the slope is 5. PART B: Which function has a greater y-intercept? Answer: g(x) has a greater y-intercept because it is positive 1 and the y-intercept from the table is -1.

OpenStudy (anonymous):

x f(x) −1 −11 0 −1 1 9 g(x) g(x) = 5x + 1

OpenStudy (anonymous):

x f(x) -1 -11 0 -1 1 9

OpenStudy (welshfella):

sorry gotta go right now slope of f(x) is 10

OpenStudy (anonymous):

Okay, thanks for everything!